you are looking to retire in 42 years and plan on making monthly payments into your account to save for retirement. you believe that you will need 3.25 million in your account when you retire. you plan on investing your money back into the stock market and know from your class at the university of toledo that you should expect an average return of 10% per year. How much do you need to save each month
Answer #1

Monthly expected return (rate)=10%/12=0.833%, tenure(nper)=42*12=504, Future Value=3,250,000, PV=0, PMT=?

L M rate nper FV PV PMT K 0.833% 504 3,250,000 =PMT(K1, K2,K4,K3,0) PMT(rate, nper, pv, [fv], [type])

rate nper FV PV PMT 0.833% 504 3,250,000 ($419.66)

Hence, you need to save $419.66 each month to reach your goal.